### v1

@@ 2.5 Evaluation metrics
where $y_{t}$ and $\hat{y}_{t}$ represent the actual and predicted room temperatures at time $t$, across $T$ time steps.
@@ 3.1 Dataset
The main subject of investigation is the inhabited experimental unit Urban Mining and Recycling (UMAR), located at the Swiss Federal Laboratories for Materials Science and Technology (Empa) in Dübendorf [32], as shown in Figure 3. This unit comprises several rooms equipped with various indoor and outdoor sensors, which have been recording data over several years. Our study concentrates on the five most frequently used rooms: two bedrooms (R272, R274), a living room (R273), and two bathrooms (R275, R276). The utility room is not regarded for our study. A comprehensive overview of all weather, building, and room sensors utilized in this study is provided in Table 2. The dataset for our case study is recorder at a 1-minute resolution for all sensor variables spanning the two years 2020 and 2021. For the whole dataset, the percentage of missing values is below 1%. After linearly interpolating the missing values, the dataset is aggregated to 15-minutes resolution. The dataset is split into a 50%/50% training/test set, with the year 2020 used for training and the year 2021 used for testing.
@@ 3.3 Model architecture
For the data-driven part, we compare three models of differing complexity: LR, FFNN and RF. While LR and RF implementations are used from Scikit-learn [31], the FFNN is implemented in Pytorch [30]. All three models are trained on standardized feature inputs. The LR model used is a multiple ordinary least squares regression with an intercept. The FFNN consists of two layers with 128 neurons activated by the sigmoid function. It is trained using a batch size of 32, with a maximum of 1000 epochs, early stopping with a patience of 10, and a validation split of 20%. The model is optimized using the Adam optimizer and the mean squared error loss function. In case of the RF, the number of forest trees is 300, the splitting criterion squared error, the minimum samples split 2 and minimum samples leaf 1. The specific FFNN and RF architectures are determined through hyperparameter grid search for the data-driven case. The chosen hyperarameters are maintained consistent across all hybrid approaches as well as all documentation/sensor scenario in order to isolate the effects of each experimental condition.
@@ 4.1 Prediction performance
In order to comprehensively evaluate the the hybrid models’ prediction performance in real-world settings, we explore their performance in three different documentation/sensor scenarios. The results for the three documentation/sensor scenarios employing the time series regression framework are illustrated in Figure 4, presented as a MAPE boxplot across all rooms. Across all methods, the average MAPE (indicated by the green triangles) decreases as the level of documentation and sensor data increases from W-scenario to WBR-scenario. Notably, for the best performing method, Residual-FFNN, the MAPE improves by 0.88%, decreasing from 4.55 % in the W-scenario to 3.67 % in the WBR-scenario. This indicates that while the weather features and archetypal building characteristics provide a robust foundation for prediction, significant gains are achieved through the inclusion of more more sensors and documentation data. Furthermore, all hybrid methods,except for the surrogate models, outperform the EP simulator in terms of MAPE. Across all scenarios, the hybrid models exhibit performance comparable to the purely data-driven LR and FFNN. Nevertheless, we can see a slight performance advantage for the Residual-FFNN in the WBR-scenario.
Figure 4: MAPE boxplot for the three scenarios W, WB and WBR. The green triangle and line denote mean resp. median of the error distribution.
Figure 5 presents the performance of the WBR scenario, broken down by room. While the Residual-FFNN demonstrates the best overall performance (indicated by the grey-dotted line), there are notable variations across different room types. Specifically, the Residual-FFNN achieves the lowest MAPE in bedroom 272, whereas the Residual-RF performs better in the living room 273 and bedroom 274. For the two bathrooms, the Augmentation-FFNN yields the lowest MAPE for bathroom 275 and the data-driven RF the best MAPE performance. An example of the Residual-FFNN’s superior performance is highlighted in Figure 6. During extended window openings from March 4th to 7th and March 11th to 15th (shaded in grey), the Residual-FFNN accurately captures the extreme temperature fluctuations by leveraging the EP model. In contrast, the other hybrid approaches fail to predict this behavior effectively. A similar effect can be observed in the residual approaches of LR and RF.
@@ 4.2 Explainability within model
Figure 10 displays the hierarchical Shapley dendrogram for the inputs of the best-performing method, Residual-FFNN. With a few minor exceptions, four distinct groups can be identified. The orange cluster is composed mainly of outdoor temperature and room mass flow variables. The green cluster includes wind, window, and occupancy variables. The red and violet clusters are dominated by EP simulated inputs, solar radiation, and relative humidity, while the brown group consists of setpoints and shading variables. Notably, Table  3 exposes that the red cluster holds the highest relative importance among all clusters for room 273. This indicates that solar radiation, simulated air temperatures and relative humidity are the most impactful input group. The strong influence of solar radiation is consistent with the observation that the living room has three large windows. At the same time, the importance of the simulated air temperatures can be attributed to the nature of the residual model. If we cut the dendogram at the second level, the red, violet and brown groups together have an aggregated hierarchical Shapley value that is more than double that of the combined orange and green clusters for room 273. We can interpret features directly related to indoor temperature (simulated air temperatures and temperature set points) as well as irradiation related features (solar radiation and shading) contribute significantly more on average than outside temperature, wind features, window openings and occupancy. A similar effect is observed for the two bedrooms, except that the violet group has gained importance due to the simulated temperature of the respective bedrooms being in this group. In case of the bathrooms, we notice a decrease of importance in all groups besides the red cluster, of which the simulated temperatures of bathrooms are part of. Such an effect can be attributed to the more isolated dynamics of the bathrooms having no windows.
Figure 10: Dendogram for input of Residual-FFNN in the WBR-scenario.
Group 1 Goup 2 Group 3 Group 4 Group 5
@@
Table 3: Hierarchical Shapley values for all dendogram groups in the case of the Residual-FFNN.
@@ 4.3 Explainability comparison between models
In order to compare the influences of the input cluster groups across the different hybrid approaches, Figure  11 shows the averaged absolute hierarchical Shapley values across rooms for LR, FFNN and RF. Note that the cluster groups of the assistant and residual approaches follow the same form shown in Figure 10 due to the additional simulated input, and are henceforth denoted as AR-clustering. The data-driven, augmentation and surrogate approach follow the same form given in Figure 16 having no additional simulated input and are denoted as DAS-clustering. In princile, the DAS-clustering resembles the AR-clustering for most parts, except for the absence of simulated inputs, the violet group being dissolved into the orange and green one as well as two set point input being merged into the green cluster. Moreover, note that the values between LR, FFNN, and RF hybrid approaches should not be compared directly but only in relative terms, since the Shapley value calculation is highly dependent on the model’s inner workings. The most informative observation in Figure  11 is the high importance of the red cluster relative to the other clusters in case of the assistant and residual approaches for RL, FFNN and RF. In contrast, the DAS-clustering bar plots show a more balanced picture with the orange group having the highest importance. This suggests that the assistant and residual approach are provided with most parts of the dynamics through the simulated inputs in the red group, while the data-driven, augmentation and surrogate approach have to learn the temperature dynamics primarily through the sensor data.
@@
Table 4: Model native importance measures contrasted with hierarchical Shapley values for Residual-RL, -FFNN and -RF
@@ 4.4 Impact of data quantity on performance
In practice, building sensor data may not abundantly available due to cost and infrastructure challenges or privacy concerns. In this light, we investigate the data quantity dependence of the hybrid models with limited data. In these experiments, we reduce the amount of training data and evaluate the hybrid models on the same test set (year 2021) as before. The MAPE results for these experiments, using limited training data, are displayed in Figure 12. The models are trained with a progressively decreasing number of months, ranging from January 15th to July 15th. The results for 12 months of training data are shown as a reference on the left side of the figure. For all experiments, we keep the same calibration of the EnergyPlus model as in the 12-months set-up, assuming ideal conditions.
@@
Among the models, the Residual-FFNN performs best, while the Surrogate-FFNN consistently shows the highest MAPE across all training data sizes. Additionally, the performance gap between the FFNN and Residual-FFNN widens as the amount of training data increases. Notably, from three months of training data, the limited data significantly impacts the FFNN’s performance. In contrast, the Residual-FFNN effectively leverages the physical information from the EP simulation to generate more accurate predictions. An illustration of this can be seen in Figure 13, which shows predictions over the entire test set using only 1 month of training data. Similar effects can be observed in cases of LR and RF, depicted in Appendix Figure 19 and Figure 13. In summary, we see that the performance declines with decreasing training data while the residual approach maintains the best ability to cover the temperature variations.
@@ 5 Discussion
Our results indicate that more detailed documentation for the physics-based model and increased sensor availability significantly improve prediction accuracy. In scenarios with limited documentation and sensor data, hybrid approaches do not show substantial improvements over purely data-driven models. However, in scenarios with comprehensive documentation and abundant sensor data, hybrid methods demonstrate considerable performance enhancements. This suggests that hybrid approaches reach their full potential only when sufficient sensor measurements related to temperature dynamics are available. Additionally, we observe that the most influential inputs vary depending on the hybrid approach used. For hybrid methods incorporating physics-based simulations as additional inputs (assistant and residual approaches), the feature group containing simulation data exhibits the highest relevance, indicating that the physics-based simulations provide the majority of information about temperature dynamics. In contrast, for hybrid approaches like augmentation and surrogate, the importance is more evenly distributed across feature groups, indicating that these methods learn the dynamics directly from sensor data.
@@ 6 Conclusion
In this study, we assess the performance of four predominant hybrid approaches along with purely data-driven and physics-based models for predicting building temperatures. To mimic challenging real-world conditions, we evaluate these hybrid methods across three common scenarios that vary in documentation detail and sensor availability. Furthermore, we explore the explainability of the hybrid approaches with the hierarhical Shapley value to account for feature dependencies. The results show that the residual approach is the most effective hybrid approach for temperature prediction, while the surrogate approach consistently exhibits the lowest accuracy. In terms of explainability, we showcased the adequacy of the hierarchical Shapley value to interpret the hybrid approaches. Future development of our study could involve conducting experiments on a larger scale, including a greater number of buildings and various building types beyond residential. In addition, while we considered multiple data scenarios, further investigation into data availability could involve more granular levels of building documentation. Another area for improvement is the assumption of ideal physics-based calibration when assessing the impact of limited training data on performance. We plan to address this by re-calibrating the physics-based model based on available data.
@@ Appendix A Additional results
Figure 16: Dendogram for input of datat-driven, augmentation, and surrogate approach in the WBR-scenario.
@@
Figure 21: Bar plot of the mean absolute linear effects and hierarchical Shapley values for all features.
Figure 22: Bar plot of the mean absolute sensitivity coefficients and hierarchical Shapley values for all features.
Figure 23: Bar plot of the mean impurity decrease and hierarchical Shapley values for all features.

### v2

@@ 2.5 Evaluation metrics
where $y_{t}$ and $\hat{y}_{t}$ represent the real and predicted room temperatures at time $t$, across $T$ time steps.
@@ 3.1 Dataset
The main subject of investigation is the inhabited experimental unit Urban Mining and Recycling (UMAR), located at the Swiss Federal Laboratories for Materials Science and Technology (Empa) in Dübendorf [47], as shown in Figure 3. In terms of representativeness, previous research indicates that the UMAR unit is representative of the average Swiss residential building stock around 2050 [48]. The unit comprises several rooms equipped with various indoor and outdoor sensors, which have been recording data over several years. Our study concentrates on the five most frequently used rooms: two bedrooms (R272, R274), a living room (R273), and two bathrooms (R275, R276). The utility room is not regarded for our study. A comprehensive overview of all weather, building, and room sensors utilized in this study is provided in Table 2. The dataset for our case study is recorder at a 1-minute resolution for all sensor variables spanning the two years 2020 and 2021. For the whole dataset, the percentage of missing values is below 1%. After linearly interpolating the missing values, the dataset is aggregated to 15-minutes resolution. The dataset is split into a 50%/50% training/test set, with the year 2020 used for training and the year 2021 used for testing.
@@ 3.3 Model architecture
For the data-driven part, we compare three models of differing complexity: LR, FFNN and RF. While LR and RF implementations are used from Scikit-learn [49], the FFNN is implemented in Pytorch [50]. All three models are trained on standardized feature inputs. The LR model used is a multiple ordinary least squares regression with an intercept. The FFNN consists of two layers with 128 neurons activated by the sigmoid function. It is trained using a batch size of 32, with a maximum of 1000 epochs, early stopping with a patience of 10, and a validation split of 20%. The model is optimized using the Adam optimizer and the mean squared error loss function. In case of the RF, the number of forest trees is 300, the splitting criterion squared error, the minimum samples split 2 and minimum samples leaf 1. The specific FFNN and RF architectures are determined through hyperparameter grid search for the data-driven case. The chosen hyperparameters are maintained consistent across all hybrid approaches as well as all documentation/sensor scenario in order to isolate the effects of each experimental condition.
@@ 4.1 Prediction performance
In order to comprehensively evaluate the hybrid models’ prediction performance in real-world settings, we explore their performance in three different documentation/sensor scenarios. The results for the three documentation/sensor scenarios employing the time series regression framework are illustrated in Figure 4, presented as a MAPE boxplot across all rooms. Across all methods, the average MAPE (indicated by the green triangles) decreases as the level of documentation and sensor data increases from W-scenario to WBR-scenario. Notably, for the best performing method, Residual-FFNN, the MAPE improves by 0.88%, decreasing from 4.55 % in the W-scenario to 3.67 % in the WBR-scenario. This indicates that while the weather features and archetypal building characteristics provide a robust foundation for prediction, significant gains are achieved through the inclusion of more more sensors and documentation data. Furthermore, all hybrid methods,except for the surrogate models, outperform the EP simulator in terms of MAPE. Across all scenarios, the hybrid models exhibit performance comparable to the purely data-driven LR and FFNN. Nevertheless, we can see a slight performance advantage for the Residual-FFNN in the WBR-scenario.
Figure 4: MAPE boxplot for the three scenarios W, WB and WBR between real indoor temperature and model prediction. The green triangle and line denote mean resp. median of the error distribution.
@@
Figure 5 presents the performance of the WBR scenario, broken down by room. While the Residual-FFNN demonstrates the best overall performance (indicated by the grey-dotted line), there are notable variations across different room types. Specifically, the Residual-FFNN achieves the lowest MAPE in bedroom 274, whereas the Residual-RF performs better in the living room 273 and bedroom 272. For the two bathrooms, the Augmentation-FFNN yields the lowest MAPE for bathroom 276 and the data-driven RF the best MAPE performance for bathroom 275. An example of the Residual-FFNN’s superior performance is highlighted in Figure 6. During extended window openings from March 4th to 7th and March 11th to 15th (shaded in grey), the Residual-FFNN accurately captures the extreme temperature fluctuations by leveraging the EP model. In contrast, the purely data-driven model and other hybrid approaches fail to predict this behaviour effectively. The FFNN shows difficulties to learn relationships that generalize to such extreme window behaviour. The Assistant-FFNN approach seems unable to beneficially use the EP simulation as feature, even deteriorating performance compared to FFNN in these cases. In the augmentation approach, we observe a slight improvement over FFNN, indicating that the two-step learning approach can bring advantages in such situations. A similar effect can be observed in the residual approaches of LR and RF Figure 16 and Figure  17. The mean performance of all methods for MAPE, MAE and RMSE is given in Figure 5.
@@
4.2 Influence of historical sensor data
The influence of historical sensor data can give insights into the impact of short- and long-term effects on prediction performance. To this end, we examine a Long Short-term Memory (LSTM) Network [51]. To be comparable to the FFNN, we employ a two layer network with 128 neurons each and train it in the same fashion.
In Figure 8, we display the MAPE for the hybrid approaches with FFNN and LSTM. We observe that the Residual approach shows the best overall performance, with a slight advantage of Residual-FFNN over Residual-LSTM. This could indicate that the residual approach has a superior ability to leverage short-term historical effects. Nevertheless, when looking at the room-wise performance, we see that the pure data-driven models outperform the residual approach for the two bathrooms. This phenomenon could be traced back again to the peculiar dynamics of the bathrooms, complicating their physics-based modeling.
Figure 8: Comparison between hybrid approaches of FFNN and LSTM for the case of one hour history integration.
Figure 9 shows the MAPE performance of the Residual-FFNN and Residual-LSTM for increasing hours of history. In order to isolate the effect of history integration, the model architectures are kept constant. We observe that the lowest error is achieved by the Residual-FFNN with 1 hour of historical sensor data. Interestingly, the Residual-LSTM also displays its best performance with 1 hour history integration. This could indicate that short-term effects in UMAR such as solar radiation are predominant, and integrating longer-term effects may not bring additional benefits. Further, we observe a large performance degradation with increasing hours for Residual-FFNN. This suggests an underfitting of the model, since one additional hour of history corresponds to an addition of about 200 features. Conversely, the Residual-LSTM has a superior ability to integrate the historical data having a substantially lower error for increasing hours.
Figure 9: Comparison between Residual-FFNN and Residual-LSTM when including different hours of sensor history. Note that 0 hours of history represents the direct input-output mapping from features to target at the same time step.
@@ 4.3 Explainability within model
Figure 12 displays the hierarchical Shapley dendrogram for the inputs of the best-performing method, Residual-FFNN. With a few minor exceptions, four distinct groups can be identified. The orange cluster is composed mainly of outdoor temperature and room mass flow variables. The green cluster includes wind, window, and occupancy variables. The red and violet clusters are dominated by EP simulated inputs, solar radiation, and relative humidity, while the brown group consists of setpoints and shading variables. Notably, Table  3 exposes that the red cluster holds the highest relative importance among all clusters for room 273. This indicates that solar radiation, simulated air temperatures and relative humidity are the most impactful input group. The strong influence of solar radiation is consistent with the observation that the living room has three large windows. At the same time, the importance of the simulated air temperatures can be attributed to the nature of the residual model. If we cut the dendrogram at the second level, the red, violet and brown groups together have an aggregated hierarchical Shapley value that is more than double that of the combined orange and green clusters for room 273. We can interpret features directly related to indoor temperature (simulated air temperatures and temperature set points) as well as irradiation related features (solar radiation and shading) contribute significantly more on average than outside temperature, wind features, window openings and occupancy. A similar effect is observed for the two bedrooms, except that the violet group has gained importance due to the simulated temperature of the respective bedrooms being in this group. In case of the bathrooms, we notice a decrease of importance in all groups besides the red cluster, of which the simulated temperatures of bathrooms are part of. Such an effect can be attributed to the more isolated dynamics of the bathrooms having no windows.
Figure 12: Dendrogram for input of Residual-FFNN in the WBR-scenario.
Group 1 Group 2 Group 3 Group 4 Group 5
@@
Table 3: Hierarchical Shapley values for all dendrogram groups in the case of the Residual-FFNN.
@@ 4.4 Explainability comparison between models
In order to compare the influences of the input cluster groups across the different hybrid approaches, Figure  13 shows the averaged absolute hierarchical Shapley values across rooms for LR, FFNN and RF. Note that the cluster groups of the assistant and residual approaches follow the same form shown in Figure 12 due to the additional simulated input, and are henceforth denoted as Assisted/Residual-clustering (AR-clustering). The data-driven, augmentation and surrogate approach follow the same form given in Figure 18 having no additional simulated input and are denoted as Data-driven/Augmentation/Surrogate-clustering (DAS-clustering). In principle, the DAS-clustering resembles the AR-clustering for most parts, except for the absence of simulated inputs, the violet group being dissolved into the orange and green one as well as two set point input being merged into the green cluster. Moreover, note that the values between LR, FFNN, and RF hybrid approaches should not be compared directly but only in relative terms, since the Shapley value calculation is highly dependent on the model’s inner workings. The most informative observation in Figure  13 is the high importance of the red cluster relative to the other clusters in case of the assistant and residual approaches for RL, FFNN and RF. In contrast, the DAS-clustering bar plots show a more balanced picture with the orange group having the highest importance. This suggests that the assistant and residual approach are provided with most parts of the dynamics through the simulated inputs in the red group, while the data-driven, augmentation and surrogate approach have to learn the thermodynamics primarily through the sensor data.
@@
Table 4: Model native importance measures contrasted with hierarchical Shapley values for Residual-LR, -FFNN and -RF
@@ 4.5 Impact of data quantity on performance
In practice, building sensor data may not be abundantly available due to recent sensor installation, temporary sensor deployment or sensor failures. In this light, we investigate the data quantity dependence of the hybrid models with limited data. In these experiments, we reduce the amount of training data and evaluate the hybrid models on the same test set (year 2021) as before. The MAPE results for these experiments, using limited training data, are displayed in Figure 14. The models are trained with a progressively decreasing number of months, ranging from January 15th to July 15th. The results for 12 months of training data are shown as a reference on the left side of the figure. For all experiments, we keep the same calibration of the EnergyPlus model as in the 12-months set-up, assuming ideal conditions.
@@
Among the models, the Residual-FFNN performs best, while the Surrogate-FFNN consistently shows the highest MAPE across all training data sizes. Additionally, the performance gap between the FFNN and Residual-FFNN widens as the amount of training data increases. Notably, from three months of training data, the limited data significantly impacts the FFNN’s performance. In contrast, the Residual-FFNN effectively leverages the physical information from the EP simulation to generate more accurate predictions. An illustration of this can be seen in Figure 15, which shows predictions over the entire test set using only 1 month of training data. Similar effects can be observed in cases of LR and RF, depicted in Appendix Figure 19, Figure 20, Figure 21 and Figure 22. In summary, we see that the performance declines with decreasing training data while the residual approach maintains the best ability to cover the temperature variations.
@@ 5 Discussion
Our results indicate that more detailed documentation for the physics-based model and increased sensor availability significantly improve prediction accuracy. In scenarios with limited documentation and sensor data, hybrid approaches do not show substantial improvements over purely data-driven models. However, in scenarios with comprehensive documentation and abundant sensor data, hybrid methods demonstrate considerable performance enhancements. This suggests that hybrid approaches reach their full potential only when sufficient sensor measurements related to thermodynamics are available. Additionally, we observe that the most influential inputs vary depending on the hybrid approach used. For hybrid methods incorporating physics-based simulations as additional inputs (assistant and residual approaches), the feature group containing simulation data exhibits the highest relevance, indicating that the physics-based simulations provide the majority of information about thermodynamics. In contrast, for hybrid approaches like augmentation and surrogate, the importance is more evenly distributed across feature groups, indicating that these methods learn the dynamics directly from sensor data.
@@
A limiting factor in the analysis of the limitations of training data is the assumption of ideal calibration of the physics-based model. With limited data available for calibration, the process may be suboptimal, leading to degraded performance of the physics-based model. We hypothesize that a residual-based approach could compensate for such systematic errors, whereas other hybrid approaches, such as surrogate models, may experience more significant performance limitations. Further studies are required to investigate the effects of calibration under these constrained conditions.
@@
For real-world deployment in control applications, the primary challenges are expected to include handling sensor faults, managing missing data, and adapting to usage behaviour changes due to shifting occupancy patterns. Compared to a purely physics-based model, the hybrid model is anticipated to have a minimal computational burden at inference. In selecting a suitable hybrid model, various criteria must be considered. The residual approach is best suited for compensating systematic biases in the physics-based model, such as deviations resulting from calibration errors or weather variability. The assisted approach similarly incorporates the physics-based simulation as feature, but its reliance on data-driven training means that the contribution of the physics-based simulation is determined by the model itself, offering practitioners less control. The surrogate model, in contrast, is primarily constrained by the performance of the underlying physics-based model making it less suitable for similar operation settings. The augmentation approach refines the surrogate model through fine-tuning on real data, potentially improving accuracy. However, finding an optimal tuning strategy can be complex and resource-intensive.
@@ 6 Conclusion
In this study, we assess the performance of four predominant hybrid approaches along with purely data-driven and physics-based models for predicting building temperatures. To mimic challenging real-world conditions, we evaluate these hybrid methods across three common scenarios that vary in documentation detail and sensor availability. Furthermore, we explore the explainability of the hybrid approaches with the hierarchical Shapley value to account for feature dependencies. The results show that the residual approach is the most effective hybrid approach for temperature prediction, while the surrogate approach consistently exhibits the lowest accuracy. In terms of explainability, we showcased the adequacy of the hierarchical Shapley value to interpret the hybrid approaches. While the residual approach proved to be most effective in our settings, further improvements could involve weighting samples based on a certainty metric for correction. This would enable the residual model to prioritize time periods where higher error corrections are more significant. Future development of our study could involve conducting experiments on a larger scale, including a greater number of buildings and various building types beyond residential. In addition, while we considered multiple data scenarios, further investigation into data availability could involve more granular levels of building documentation. Another area for improvement is the assumption of ideal physics-based calibration when assessing the impact of limited training data on performance. We plan to address this by re-calibrating the physics-based model based on available data. An alternative research direction could explore an online setting for hybrid approaches, incorporating frequent re-calibration of the physics-based model as well as continuous re-training of the data-driven model.
@@
Data availability
The code and dataset are publicly available at the GitHub repository.
Acknowledgement
We thank all involved members from the Urban Energy Systems Laboratory at Empa and Intelligent Maintenance and Operations Systems Laboratory at EPFL. Special thanks go to Fazel Khayatian for insightful discussions about Building Energy Modeling and the nestli platform. This project is funded by Empa research & development grant 5213.00276.
@@ Appendix A Additional results
Figure 18: Dendrogram for input of data-driven, augmentation, and surrogate approach in the WBR-scenario.
@@
Figure 23: Bar plot of the mean absolute linear effects and hierarchical Shapley values for all features in the case of Residual-LR.
Figure 24: Bar plot of the mean absolute sensitivity coefficients and hierarchical Shapley values for all features in the case of Residual-FFNN.
Figure 25: Bar plot of the mean impurity decrease and hierarchical Shapley values for all features in the case of Residual-RF.